These are the news starting 01 Jan 2005
Click here for the news of the previous years.
|28 Dec 05||Again a serial comms problem starting the 26th. Ok since midnight today. Previous days graphs will be rebuild. No data loss.|
|18 Dec 05||Uploaded nov05.dat file in data archive.|
|17 Dec 05||Yesterday (16Dec05) again comm's fault between logger and computer. Cleaned all contacts and lowered transmission speed to from 9600 to 4800... Let's wait and see.|
|11 Dec 05||A problem with the main computers. Logger halted and everything restarted. Previous days graphs will be rebuilt during the next days.|
|25 Nov 05||NOx sensor working normally again|
|22 Nov 05||Changed membranes of main NOx sensor pump. Sensor now working ok again, but several parameters like pressure and flow are close to limit.|
|21 Nov 05||
1. Maintenance work on all gas sensors. There is
a problem with the membrane-pump of the NOx sensor; NO and NO2
data may be dubious.
2. Datalogger replaced by spare; hopefully the spurious communication errors will disappear...
|16 Nov 05||Updated version of presentation Gloabl Warming? in Papers archive. Please read using IE as the conversion from ppt to html gives poor results with other browsers.|
|05 Nov 05||Again a communication problem between logger and computer. This could be an serial driver chip going bad... the problem is intermittent, and so difficult to trace. Restarted today at 12:00.|
|01 Nov 05||Upload of oct05.dat file|
|29 Oct 05||Communication breakdown between logger and main computer. No data lost, just cosmetic problem with the graphs!|
|15 Oct 05||Main computer failed again today at 04:30; has been replaced at 15:30. No data lost, but Ascii file of today's data has several multiple identical lines due to resync.. This does not hurt!|
|06 Oct 05||
Inspection of August data file showed that the
main sample pump broke down at 29Aug05 09:30. The aug05.dat file
has been corrected for this (see metadata in first lines).
Upload of sep05.dat file
|05Oct 05||Main logger computer went erratic starting 02/10/05 at 22:30. This is now temperarly solved; possibly this computer must be replaced. No data have been lost, due to internal logger storage.|
|14 Sep 05||Uploaded aug05.dat file|
|13 Sep 05||Ground ozone sensor is working correctly again.|
|12 Sep 05||Main sample pump is broken...a kludge has been implemented for the ground ozone sensor sample inlet; this sensor is working again. New pump has been ordered, and will be received during the next days.|
|2 Sep 05||Ground ozone sensor malfunction; no repair possible during next 2 weeks. Sorry!|
|24 Aug 05||Uploaded jul05.dat file|
|23 Aug 05||
site has been accessed more than 100000 times since it's start
Big thanks to all our visitors and contributors!
|30 Jul 05||
1. the total dobson page has not been correctly updated during the last 2 weeks, due to a software problem. Should be ok by now
2. the rain sensor was blocked during the last 2 days: this has been repaired, and the missing data imported into 7days.dat and all.dat from our Davis backup station (29/07 rainfall was missing)
|13 Jul 05||Added link to ground ozone data from Trier (D) (about 35 km distance) in live data page.|
|05 Jul 05||Uploaded jun05.dat file.|
|29 Jun 05||Ozone sensor problem since tomorrow evening: have been disturbed during maintenance work and left the O3 sensor in zero-test mode. Sorry.... Normal operation resumed at 17:00 UTC|
|14 Jun 05||There was a power-supply problem with the meteo01 server (which holds the lightning screens and works as a mirror for meteo.lcd.lu). Power-supply replaced, and new image made today. There may be some disruption in the lightning sequences of yesterday and today.|
|10 Jun 05||Uploaded file with total ozone measurements from 23 April 2004 to 10 June 2005 (instrument #5375); original Microtops format and Excel format.|
|02 Jun 05||Uploaded April05 and May05 data into archive. Sorry for the delay with April, I was too busy working on our exhibition of vintage computers and calculators!|
|01Jun 05||Recalibration of gaz sensors, maintenance work.|
|30 Apr 05||Recalibration of gaz sensors, maintenance work.|
|05 Apr 05||Datalogger memory overflow problem. Solved at 19:00 UTC with a download of all data and a logger restart.|
|22 Mar 05||Recalibration and maintenance work on O3, NOx and CO2 sensors from 16:00 to 18:00|
|16 Mar 05||Restarted WBGT sensor by enabling wet bulb sensor at 09:00 (wet bulb was dry during the winter to avoid freezing damage)|
|05 Mar 05||
1. Some recalibration needed after Envitec's maintenance (CO2
and NOx). The O3 readings are definitively back to normal.
2. Added UVA graphs on main page of near-live data (today_01.html); this required a slight reshuffling of the icons to make room for the new graphs. I tried to change as little as feasible. UVA irradiance is measured in Watt per m2 [W/m2]; this type of radiation is known to present a much smaller skin-cancer risk as does UVB, but if used excessively, will speed up aging of the skin by weakening collagen tissue. The UVA data were always present in the data files, but have not been plotted in the past.
|04 Mar 05||Maintenance by Envitec. Cleaning of part of the teflon tubing of the inlet to the O3 sensor. Extremely high O3 values after that, than back to normal. Reason is unknown.|
|03 Mar 05||Uploaded feb05.dat file|
|19 Feb 05||
1. Uploaded jan05.dat file
2. Reinstalled UVA SolarLight sensor, which is back from calibration. There was no need to change any factor; this instrument is quite stable!
3. Sorry for not yet having added the 2004 data to the big Excel file; I'm too busy preparing an exhibition of vintage computers and calculators...so this has to wait a bit. If you really are in need, please drop me an email!
|27 Jan 05||Maintenance work and recalibration of all gas sensors.|
|13 Jan 05||Recalibration of the CO2 sensor.|
|10 Jan 05||The original (and unreliable) Bittsens software for reading the radioactivity sensor has been replaced by a "home-grown" version written by Claude Baumann (as a Labview VI). The PIC microprocessor converting the readings into voltage has also been reprogrammed. Thanks Claude for this good work!|
|09 Jan 05||
1. uploaded Dec. 2004 data files
2. uploaded full year data file (= concatenated monthly files) as 2004.zip
3. the 2004 data have not yet been integrated into the BIG Excel file... please be patient!
|04 Jan 05||The UVA sensor (ch.10) has been dismounted to be sent back to Solar Light Co. for recalibration; please use sensor UVA_DT (channel 7) as a replacement for UVA readings.|
|03 Jan 05||
Repair work at 14:30:
1. Mounted new pump on the ozone sensor (ch.14); usual API sensor again online.
2. Replaced power-supply of CO2 sensor by a new one; recalibration of sensor. Check for jump in signal: previous readings may have been too high.
|02 Jan 05||The usual New Year problem with the 7 days graph (the logger data do not contain the year information, so the automatically graphing by Gnuplot gets confused). Restarted 7days graphs at 01Jan2005 06:00|